OpenStudy (anonymous):

The perimeter of an equilateral triangle is 24, what is the length of the altitude of this triangle

OpenStudy (anonymous):

The altitude of an equilateral triangle will bisect the side it intersects. When, that happens there will be a short leg that equals 12 and a hypotenuse that equals 24. The altitude is the long leg, so its length is 12sqrt(3)

OpenStudy (anonymous):

I made a mistake, the hypotenuse of the triangle is 8 because the length of a side is 8. Therefore, the altitude is 4sqrt(3)

OpenStudy (anonymous):

24/3=8 , that is each side of the triangle.then your base will be 8. take 1/2 of the base. from that point take a line to the tip of the triangle , that is the altitude.now, the sloped side is 8 , the base is 4. now use the formula a squared + b squared = c squared. a^2+b^2 = c^2 , a^2 + 16 = 64 , a^2 = 64 - 16 , a^2 = 48 ,square root of a^2 = square root of 48 = a = 6.928

OpenStudy (anonymous):

to explain what brangler did besides want to confuse you and make you think he's a genius was use a trig formula which says the tan. of angle A(60 degrees)is = o( a,the altitude side) / 4 , (the adjacent side) . the tangent of 60 degrees is 1.732050808.therefore 1.732= 0 or altitude or a / 4 , ( the adjacent side) . work the formula you get the same answer. ps don't have to be a genius to work any math problem. tip : when explaining a math problem , try and not leave out any steps and explain it so the person can understand it or please don't waste everyones time. thanks